# Rua Dr. Bento Teobaldo Ferraz 271, 01140-070, São Paulo, SP, Brazil

### v2

@@ 4.2 Lifting from three-point functions
With explicit $(h,\bar{h})=(J,0)$ primaries in hand, our goal is to show that none of them remain conserved in the IR for $N>4$. In other words, the anomalous dimension matrix for each $J$ should have all non-zero eigenvalues at some loop order. This order is guaranteed not to be $O(g_{\sigma})$ because of chirality. It is also guaranteed not to be $O(g_{\epsilon}^{n})$ for any $n$ since the pure $\epsilon$ deformation leaves minimal models decoupled. This leaves $O(g_{\sigma}^{2})$ as the next simplest possibility which turns out to be the correct one. Once it is established that currents lift at this order, higher orders will not be able to restore conservation since we are working with asymptotically large $m$.
The recipe of conformal perturbation theory tells us to show that there is no spin-$J$ current $T$ such that $T(0)\sigma(z,\bar{z})\sigma(1)T(\infty)\right>$ integrates to zero as a principal value. As advertised previously, conformal representation theory leveraged as in rt15 shows that this can be replaced with a technically simpler check — we can equivalently show that there is no spin-$J$ current $T$ such that $T(z_{1})V^{L}(z_{2},\bar{z}_{2})\sigma(z_{3},\bar{z}_{3})\right>$ vanishes for all spin-$(J-1)$ divergence candidates $V^{L}$. In practice, the algorithm described above produces $d_{J,0}$ operators $T^{K}$ and more than $d^{\prime}_{J,1}$ operators $V^{L}$ because duplicates are slow to remove at finite $N$. From this data, we can proceed to compute the matrix
@@
and stop as soon as it becomes impossible to find a vanishing linear combination of the rows. In favourable cases, one can conclude this by computing only $d_{J,0}$ out of the more than $d^{\prime}_{J,1}$ columns and seeing that this is already enough to give the matrix full rank. [^14] Clearly, it would be even more favourable to also sort the operators involved into $S_{N}$ irreps since these give (54) a block structure. Making this parallel is non-trivial because, even though the $N$ copy three-point functions are all different, they share a dependence on many of the same single-copy two-point and three-point functions. These building blocks should be memoized in a way that makes them accessible to each process.
@@
The algorithm above expresses a single term of $T^{K}(z_{1})V^{L}(z_{2},\bar{z}_{2})\sigma(z_{3},\bar{z}_{3})\right>$ as a sum over permutations. This is the last step which is sensitive to either the number of copies $N$ or the index structure of the operators. From now on, we need to evaluate the sum where each term is a product of single-copy correlators. These correlators come in three types.
@@ 5.2 Other probes of enhanced symmetry
There is another sense in which $N=4$ preserves more symmetry than $N>4$ and this comes from the notion of generalized symmetry gksw14 . Symmetry actions on local operators are implemented by topological operators of codimension $1$ and in 2d these are topological defect lines TDL . If the fusion rules for topological defect lines take the form of a group law then they reproduce ordinary symmetries but more generally one must appeal to the concept of a fusion category. [^15] The data of a fusion category can be used to identify the obstructions to gauging which were previously mentioned in section 2.3. In particular, a topological defect line does not need to have an inverse. Minimal models have non-invertible symmetries implemented by Verlinde lines $\mathcal{L}_{(r,s)}$ which are labeled by the same integers as the primary operators and obey the same fusion rule. [^16] To mention the simplest example, the critical Ising model ($m=3$) has a single non-invertible Verlinde line and its action on local operators is the well known Kramer-Wannier duality.
@@
$\begin{split}\mathcal{S}_{(r^{\prime},s^{\prime})(r,s)}=(-1)^{rs^{\prime}+r^{\prime}s+1}\sqrt{\frac{8}{m(m+1)}}\sin\left(\pi rr^{\prime}\frac{m}{m+1}\right)\sin\left(\pi ss^{\prime}\frac{m+1}{m}\right).\end{split}$ (68)
@@
$\begin{split}\sin^{2}\left(\pi s^{\prime}\frac{m+1}{m}\right)=\sin^{2}\left(\pi\frac{m+1}{m}\right).\end{split}$ (69)
@@
$\begin{split}\cos\left(\pi s^{\prime}\frac{m+1}{m}\right)=\cos\left(\pi\frac{m+1}{m}\right)(-1)^{r^{\prime}-1}.\end{split}$ (70)
@@
Another point we can make for general $N$ is that when $\sigma$ is used to flow to a TQFT in the IR, it is guaranteed to have degenerate vaccua. This follows from the existence of topological defect lines which have a non-integer expectation value (empty circle) TDL . In such cases, cck24a showed that the standard statement of crossing symmetry generically fails and needs to be replaced by a modified crossing equation which takes the normalizations of different vaccua into account. A related phenomenon had previously been observed for Chern-Simons matter theories in mmpps22 . It is therefore possible that integrable or non-integrable S-matrices which are reachable from coupled minimal models can be bootstrapped along the lines of cck24b . [^17] See also Cordova:2024iti ; Cordova:2024vsq for closely related work.
@@ 6 Outlook
While our understanding of compact irrational CFTs is very limited as it is, there is a further refinement of this space of theories that is even less understood and that is theoretically even more important: the space of interacting compact irrational CFTs with a twist gap. The twist gap is defined as the infimum of $\tau=\Delta-J$ over the spectrum of Virasoro primaries. In our work, we have given credible evidence that all operators in the IR CFT have positive twist, but we have not addressed how this twist behaves at large spin. While the general expectation based on arguments of convexity is that twist should increase with spin, it is a logical possibility that the twist actually decreases with spin such that the twist gap vanishes. [^18] In fact, for theories with a $u(1)$ chiral algebra it can be proven that the $u(1)$ twist gap vanishes Benjamin:2020swg . The existence of a positive twist gap is a key assumption of many modern universal results about 2d CFTs. The lightcone modular bootstrap Collier:2016cls ; Benjamin:2019stq ; Pal:2022vqc ; Pal:2023cgk establishes the existence of families of operators whose weights approach $(c-1)/24$ and the Virasoro lightcone bootstrap Kusuki:2018wpa ; Collier:2018exn further establishes the existence of deformed double-twist trajectories. [^19] Along with trajectories of analytically continued scaling dimensions, a twist gap also influences the behaviour of averaged OPE coefficients involving heavy operators universal1 . For boundary and crosscap analogues of this result, see Kusuki:2021gpt ; universal2 ; universal3 . Isolating the leading Regge trajectory in our models and understanding its shape will be important for answering whether or not we have identified CFTs that satisfy the assumptions of these key results. This could also allow us to test these predictions, guided by how a detailed numerical understanding of the 3d Ising model led to a verification of the power and accuracy of the analytic bootstrap in $d>2$ Simmons-Duffin:2016wlq .
@@ Acknowledgements.
We are grateful for conversations with Christopher Beem, Nathan Benjamin, Jasper Jacobsen, Elias Kiritsis, Dalimil Mazáč, Hirosi Ooguri, Eric Perlmutter, Sylvain Ribault, Slava Rychkov, Adar Sharon and Yifan Wang. AA received funding from the German Research Foundation DFG under Germany’s Excellence Strategy – EXC 2121 Quantum Universe – 390833306 and is funded by the European Union (ERC, FUNBOOTS, project number 101043588). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. CB received funding from the São Paulo Research Foundation (FAPESP) grants 2019/24277-8, 2021/14335-0 and 2023/03825-2. AA thanks IPhT Saclay where part of this work was presented. CB thanks the Perimeter Institute for hospitality during the final stages of this work. Research at the Perimeter Institute is supported in part by the Government of Canada through NSERC and by the Province of Ontario through MRI.

### v3

@@ 4.2 Lifting from three-point functions
With explicit $(h,\bar{h})=(J,0)$ primaries in hand, our goal is to show that none of them remain conserved in the IR for $N>4$. In other words, the anomalous dimension matrix for each $J$ should have all non-zero eigenvalues at some loop order. This order is guaranteed not to be $O(g_{\sigma})$ because of chirality. It is also guaranteed not to be $O(g_{\epsilon}^{n})$ for any $n$ since the pure $\epsilon$ deformation leaves minimal models decoupled. This leaves $O(g_{\sigma}^{2})$ as the next simplest possibility which turns out to be the correct one. Once it is established that currents lift at this order, higher orders will not be able to restore conservation since we are working with asymptotically large $m$ and therefore small coupling.
The recipe of conformal perturbation theory tells us to show that there is no spin-$J$ current $T$ such that $T(0)\sigma(z,\bar{z})\sigma(1)T(\infty)\right>$ integrates to zero as a principal value. As advertised previously, conformal representation theory leveraged as in rt15 shows that this can be replaced with a technically simpler check — we can equivalently show that there is no spin-$J$ current $T$ such that $T(z_{1})V^{L}(z_{2},\bar{z}_{2})\sigma(z_{3},\bar{z}_{3})\right>$ vanishes for all spin-$(J-1)$ divergence candidates $V^{L}$. In practice, the algorithm described above produces $d_{J,0}$ operators $T^{K}$ and more than $d^{\prime}_{J,1}$ operators $V^{L}$ because duplicates are slow to remove at finite $N$. From these data, we can proceed to compute the matrix
@@
and stop as soon as it becomes impossible to find a vanishing linear combination of the rows. In favourable cases, one can conclude this by computing only $d_{J,0}$ out of the more than $d^{\prime}_{J,1}$ columns and seeing that this is already enough to give the matrix full rank. [^19] Clearly, it would be even more favourable to also sort the operators involved into $S_{N}$ irreps since these give (65) a block structure. We have had to skip this step because our methods are unable to ensure that all divergence candidates are linearly independent in a reasonable amount of time. Since the necessasry matrix elements do not need to be computed in any particular order, it makes sense to write parallelized code for this purpose. Nevertheless, this task is not embarrassingly parallel because, even though the $N$ copy three-point functions are all different, they share a dependence on many of the same single-copy two-point and three-point functions. These building blocks should be memoized in a way that makes them accessible to each process.
@@
The algorithm above expresses a single term of $T^{K}(z_{1})V^{L}(z_{2},\bar{z}_{2})\sigma(z_{3},\bar{z}_{3})\right>$ as a sum over permutations. We have seen that copy indices play a role in determining which permutations are absent from this sum and that the number of copies affects the coefficients for all remaining permutations. The rest of our task is insensitive to the index structure and the value of $N$ because it boils down to evaluating the terms of the sum which are products of single-copy correlators. These correlators come in three types.
@@
This concludes the list of operations that must be carried out in this approach to verify that UV currents lift. The arXiv submission for the latest version of this paper includes Python code for this purpose. It is sufficient to verify that anomalous dimensions are non-zero for the cases we have tested which are $J<10$ and $N=5,6,7$.
@@ 5.2 Other probes of enhanced symmetry
There is another sense in which $N=4$ preserves more symmetry than $N>4$ and this comes from the notion of generalized symmetry gksw14 . Symmetry actions on local operators are implemented by topological operators of codimension $1$ and in 2d these are topological defect lines TDL . If the fusion rules for topological defect lines take the form of a group law then they reproduce ordinary symmetries but more generally one must appeal to the concept of a fusion category. [^20] The data of a fusion category can be used to identify the obstructions to gauging which were previously mentioned in section 2.3. In particular, a topological defect line does not need to have an inverse. Minimal models have non-invertible symmetries implemented by Verlinde lines $\mathcal{L}_{(r,s)}$ which are labeled by the same integers as the primary operators and obey the same fusion rule. [^21] To mention the simplest example, the critical Ising model ($m=3$) has a single non-invertible Verlinde line and its action on local operators is the well known Kramers-Wannier duality.
@@
$\begin{split}\mathcal{S}_{(r^{\prime},s^{\prime})(r,s)}=(-1)^{rs^{\prime}+r^{\prime}s+1}\sqrt{\frac{8}{m(m+1)}}\sin\left(\pi rr^{\prime}\frac{m+1}{m}\right)\sin\left(\pi ss^{\prime}\frac{m}{m+1}\right).\end{split}$ (79)
@@
$\begin{split}\sin^{2}\left(\pi s^{\prime}\frac{m}{m+1}\right)=\sin^{2}\left(\pi\frac{m}{m+1}\right).\end{split}$ (80)
@@
$\begin{split}\cos\left(\pi s^{\prime}\frac{m}{m+1}\right)=\cos\left(\pi\frac{m}{m+1}\right)(-1)^{r^{\prime}-1}.\end{split}$ (81)
@@
Another point we can make for general $N$ is that when $\sigma$ is used to flow to a TQFT in the IR, it is guaranteed to have degenerate vacua. This follows from the existence of topological defect lines which have a non-integer expectation value (empty circle) TDL . In such cases, cck24a showed that the standard statement of crossing symmetry generically fails and needs to be replaced by a modified crossing equation which takes the normalizations of different vacua into account. A related phenomenon had previously been observed for Chern-Simons matter theories in mmpps22 . It is therefore possible that integrable or non-integrable S-matrices which are reachable from coupled minimal models can be bootstrapped along the lines of cck24b . [^22] See also Cordova:2024iti ; Cordova:2024vsq for closely related work.
@@ 6 Outlook
While our understanding of compact irrational CFTs is very limited as it is, there is a further refinement of this space of theories that is even less understood and that is theoretically even more important: the space of interacting compact irrational CFTs with a twist gap. The twist gap is defined as the infimum of $\tau=\Delta-J$ over the spectrum of Virasoro primaries. In our work, we have given credible evidence that all operators in the IR CFT have positive twist, but we have not addressed how this twist behaves at large spin. While the general expectation based on arguments of convexity is that twist should increase with spin, it is a logical possibility that the twist actually decreases with spin such that the twist gap vanishes. [^23] In fact, for theories with a $u(1)$ chiral algebra it can be proven that the $u(1)$ twist gap vanishes Benjamin:2020swg . The existence of a positive twist gap is a key assumption of many modern universal results about 2d CFTs. The lightcone modular bootstrap Collier:2016cls ; Benjamin:2019stq ; Pal:2022vqc ; Pal:2023cgk establishes the existence of families of operators whose weights approach $(c-1)/24$ and the Virasoro lightcone bootstrap Kusuki:2018wpa ; Collier:2018exn further establishes the existence of deformed double-twist trajectories. [^24] Along with trajectories of analytically continued scaling dimensions, a twist gap also influences the behaviour of averaged OPE coefficients involving heavy operators universal1 . For boundary and crosscap analogues of this result, see Kusuki:2021gpt ; universal2 ; universal3 . Isolating the leading Regge trajectory in our models and understanding its shape will require intensive computations of full anomalous dimension matrices but will be important for answering whether or not we have identified CFTs that satisfy the assumptions of these key results. This could also allow us to test these predictions, guided by how a detailed numerical understanding of the 3d Ising model led to a verification of the power and accuracy of the analytic bootstrap in $d>2$ Simmons-Duffin:2016wlq .
@@ Acknowledgements.
We are grateful for conversations with Christopher Beem, Nathan Benjamin, Jasper Jacobsen, Elias Kiritsis, Dalimil Mazáč, Hirosi Ooguri, Eric Perlmutter, Sylvain Ribault, Junchen Rong, Slava Rychkov, Adar Sharon, Yifan Wang and Jingxiang Wu. AA received funding from the German Research Foundation DFG under Germany’s Excellence Strategy – EXC 2121 Quantum Universe – 390833306 and is funded by the European Union (ERC, FUNBOOTS, project number 101043588). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. CB received funding from the São Paulo Research Foundation (FAPESP) grants 2019/24277-8, 2021/14335-0 and 2023/03825-2. AA thanks IPhT Saclay where part of this work was presented. CB thanks the Perimeter Institute for hospitality during the final stages of this work. Research at the Perimeter Institute is supported in part by the Government of Canada through NSERC and by the Province of Ontario through MRI.
Appendix A Further degeneracy tables
In this appendix we list irrep-refined degeneracies for $N=5,6$ using the methods of section 3.3.
$J$ $d^{(5)}_{J,0}$ $d^{(4,1)}_{J,0}$ $d^{(3,2)}_{J,0}$ $d^{(3,1,1)}_{J,0}$ $d^{(2,1,1,1)}_{J,0}$ $d^{(2,2,1)}_{J,0}$ $d^{(1,1,1,1,1)}_{J,0}$
0 1 0 0 0 0 0 0
1 0 0 0 0 0 0 0
2 0 1 0 0 0 0 0
3 0 0 0 0 0 0 0
4 1 1 1 0 0 0 0
5 0 0 0 1 0 0 0
6 2 3 2 1 0 0 0
7 0 1 1 2 0 1 0
8 4 7 5 3 0 2 0
9 1 5 4 6 2 2 0
10 7 15 13 10 1 6 0
Table 10: Degeneracies for all $N=5$ $\widehat{\mathfrak{Vir}}$ primary currents.
$J$ $d^{\prime(5)}_{J+1,1}$ $d^{\prime(4,1)}_{J+1,1}$ $d^{\prime(3,2)}_{J+1,1}$ $d^{\prime(3,1,1)}_{J+1,1}$ $d^{\prime(2,1,1,1)}_{J+1,1}$ $d^{\prime(2,2,1)}_{J+1,1}$ $d^{\prime(1,1,1,1,1)}_{J+1,1}$
0 1 1 0 0 0 0 0
1 0 1 1 1 0 0 0
2 1 2 2 1 0 1 0
3 2 5 3 4 1 1 0
4 4 9 8 7 2 5 0
5 5 16 15 16 5 9 0
6 12 31 29 28 11 19 2
7 17 53 52 57 23 37 2
8 32 93 96 98 44 72 7
9 50 160 164 182 85 126 13
Table 11: Degeneracies for all $N=5$ $\widehat{\mathfrak{Vir}}$ primary divergence candidates.
$J$ $d^{(6)}_{J,0}$ $d^{(5,1)}_{J,0}$ $d^{(4,2)}_{J,0}$ $d^{(4,1,1)}_{J,0}$ $d^{(3,1,1,1)}_{J,0}$ $d^{(3,2,1)}_{J,0}$ $d^{(3,3)}_{J,0}$ $d^{(2,2,2)}_{J,0}$
0 1 0 0 0 0 0 0 0
1 0 0 0 0 0 0 0 0
2 0 1 0 0 0 0 0 0
3 0 0 0 0 0 0 0 0
4 1 1 1 0 0 0 0 0
5 0 0 0 1 0 0 0 0
6 2 3 2 1 0 0 1 0
7 0 1 1 2 0 1 0 0
8 4 7 6 3 0 2 1 0
9 1 5 4 6 2 3 2 0
10 7 16 14 10 1 8 5 1
Table 12: List of degeneracies for all $N=6$ $\widehat{\mathfrak{Vir}}$ primary currents. Additional irreps have zero degeneracy up to spin 10.
$J$ $d^{\prime(6)}_{J+1,1}$ $d^{\prime(5,1)}_{J+1,1}$ $d^{\prime(4,2)}_{J+1,1}$ $d^{\prime(4,1,1)}_{J+1,1}$ $d^{\prime(3,1,1,1)}_{J+1,1}$ $d^{\prime(3,2,1)}_{J+1,1}$ $d^{\prime(3,3)}_{J+1,1}$ $d^{\prime(2,2,2)}_{J+1,1}$
0 1 1 1 0 0 0 0 0
1 0 1 1 1 0 1 1 0
2 1 3 3 2 0 2 1 1
3 2 6 6 6 2 5 3 0
4 5 12 15 11 4 13 6 3
5 6 22 27 26 11 28 13 5
6 15 45 57 49 24 58 25 14
7 22 78 103 100 54 119 50 25
8 44 145 199 183 106 231 92 57
9 69 255 352 349 215 438 173 102
Table 13: List of degeneracies for $N=6$ $\widehat{\mathfrak{Vir}}$ primary divergence candidates built from four $\phi_{(1,2)}$ operators. Operators in irreps with no current to recombine with are not listed.